An arrow of Mass 0.3kg is fired with a velocity of 10m/s into a wooden block if mass 0.7kg.Calculate the final kinetic energy after impact, given that the wooden block can freely move
1
Expert's answer
2019-11-29T09:10:06-0500

From the conservation of momentum:


"mv=(m+M)u"

"(0.3)(10)=(0.3+0.7)u"

"u=3\frac{m}{s}"

The final kinetic energy after impact:


"K=0.5(m+M)u^2=0.5(0.3+0.7)3^2=4.5\ J"